Cinderella has 1/4 as many bottles of magic

as her sister. In total, they had 30. How many

bottles of magic does each person have?

Define your variables and solve.


1
Expert's answer
2020-11-09T19:48:43-0500

"\displaystyle\textsf{Let the bottles of magic of Cinderella be}\, x\\


\textsf{Cinderella's sister bottle of magic is}\, x + \frac{x}{4}\\


x + x + \frac{x}{4} = 30\\
 

\frac{9x}{4} = 30\\


x = \frac{30 \times 4}{9} = \frac{40}{3} = 13.3\\

\textsf{Putting the values of}\, x,\\\textsf{we have the bottle of magic}\\\textsf{each person has.}\\


\textsf{Cinderella's sister bottle of magic is}\, \frac{40}{3} + \frac{40}{4\times 3} = \frac{50}{3} = 16.7.\\


\therefore \,\textsf{Cinderella has}\, 13.3\, \textsf{bottles of magic}\,\&\\


 \textsf{Cinderella's sister has}\, 16.7\, \textsf{bottles of magic.}"